OpenStudy (anonymous):

Solve please: \[-3\left| 4-x \right|+3\le-9\]

OpenStudy (mathmate):

−3|4−x|+3≤−9 Divide by 3 −|4−x|+1≤−3 transpose 3+1≤|4−x| 4≤|4−x| case 1: 4-x >=0, then 4-x >=4, x<=0 case 2: 4-x <0 then -(4-x)>=4 => x-4>=4 => x>=8

OpenStudy (anonymous):

Great this is an online quiz so i need fast things lol xD

OpenStudy (anonymous):

Solve please\[\left| x+7 \right|=3\]

OpenStudy (mathmate):

Same, case 1: x+7=3 => x=-4 case 2: -(x+7)=3 => x=-10